# rowhighlight linewidth=1pt, linecolor=black, innertopmargin=4pt, innerbottommargin=4pt, innerleftmargin=4pt, innerrightmargin=4pt, backgroundcolor=lightgray!15

### v1

@@ Abstract
Sparse Autoencoders (SAEs) have demonstrated significant promise in interpreting the hidden states of language models by decomposing them into interpretable latent directions. However, training SAEs at scale remains challenging, especially when large dictionary sizes are used. While decoders can leverage sparse-aware kernels for efficiency, encoders still require computationally intensive linear operations with large output dimensions. To address this, we propose KronSAE, a novel architecture that factorizes the latent representation via Kronecker product decomposition, drastically reducing memory and computational overhead. Furthermore, we introduce mAND, a differentiable activation function approximating the binary AND operation, which improves interpretability and performance in our factorized framework.
@@ 1 Introduction
Interpreting dense embeddings from large language models remains a central challenge for transparency and controllability in AI systems (Elhage et al., 2023; Heap et al., 2025). Sparse autoencoders (SAEs) have emerged as powerful tools for uncovering human-interpretable features within neural activations by enforcing activation sparsity to induce discrete-style dictionaries (Elhage et al., 2023; Gao et al., 2025; Cunningham et al., 2024). These dictionaries facilitate rigorous circuit-level semantic analysis and concept discovery, enabling fine-grained probing of model internals (Elhage et al., 2023; Cunningham et al., 2024).
However, naively scaling SAEs to the widths demanded by modern transformers leads to prohibitive compute costs—both in memory and per-step FLOPs—limiting their applicability to frontier models. Gated SAEs (Rajamanoharan et al., 2024a) address this by learning continuous sparsity masks via lightweight gating networks, achieving a Pareto improvement on the reconstruction–sparsity trade-off. Switch SAEs (Mudide et al., 2025) leverage conditional computation by routing activations among smaller expert SAEs, reducing computation by activating only a subset of experts per input. In a work Gao et al. (2025) an acceleration of TopK SAE was proposed that utilizes an optimized kernel based on efficient sparse–dense matrix multiplication. Despite these advances, the encoder remains largely unoptimized: it still performs a dense projection into the full dictionary, incurring high computational cost and limiting scalability.
In this paper, we introduce KronSAE, a novel SAE encoder that directly tackles this bottleneck. By decomposing the latent space into head-wise Kronecker factors and embedding a smooth, differentiable AND-like gating mechanism, KronSAE reduces both parameters and compute overhead while preserving reconstruction fidelity.
@@
    We identify the encoder projection as the principal scalability bottleneck in sparse autoencoders, conducting a FLOPs-aware analysis of its impact and demonstrating that targeted encoder optimizations can significantly improve computational performance while maintaining reconstruction quality.
@@
    We propose KronSAE, a Kronecker-factorised sparse autoencoder equipped with the novel $\operatorname{mAND}$ activation function. The design significantly reduces encoder cost and is compatible with existing sparse decoder kernels.
@@
    We show on multiple language models that KronSAE improves reconstruction fidelity, decreases feature absorption, and yields more interpretable latents under fixed compute.
@@ 2 Related Work
Sparse autoencoders (SAEs) enforce sparsity constraint on a latent code. Early work demonstrated that SAEs can uncover human-interpretable directions in deep models (Cunningham et al., 2024; Templeton et al., 2024), but extending them to the large latent sizes ($F\gg d$) for modern language models with large hidden dimension $d$ is expensive: each forward pass still requires a dense $\mathcal{O}(F\,d)$ encoder projection. Most prior optimization efforts focus on the decoder side. For example, Gao et al. (2025) introduce a fused sparse–dense TopK kernel that reduces wall-clock time and memory traffic, while Rajamanoharan et al. (2024a) decouple activation selection from magnitude prediction to improve the $\ell_{0}$–MSE trade-off. Separately, Rajamanoharan et al. (2024b) propose JumpReLU, a nonlinearity designed to mitigate shrinkage of large activations in sparse decoders.
Figure 1: Isoflops Pareto fronts for KronSAE vs. TopK SAE on Qwen-1.5B for different dictionary sizes $F$. 100 M tokens: KronSAE improves explained variance while heavily reducing parameters count. 500 M tokens: maintains gains across most sizes. 1000 M tokens: matches baseline quality with significantly fewer parameters; Stars denote TopK baselines and markers denote KronSAE configurations $(m,n,h)$. See details in section 4.1.
@@
Conditional-computation schemes avoid instantiating a full dictionary per token by routing inputs to a subset of expert SAEs via a lightweight gating network (Mudide et al., 2025) by employing the Mixture-of-Experts ideas (Shazeer et al., 2017), but still incur a dense per-expert encoder projection, leaving the encoder as the primary bottleneck.
Weight Factorization and Logical Activation Functions.
Outside of sparse decoders and Mixture-of-Experts, tensor-factorization methods have been used to compress large weight matrices in language models (Edalati et al., 2021). In parallel, differentiable-logic activations were introduced to approximate Boolean operators in a smooth manner (Lowe et al., 2021). Our method synthesizes these lines of work: we embed a differentiable AND-like gate into a Kronecker-factorized encoder to build rich, compositional features while preserving end-to-end differentiability.
@@ 3 Method
3.1 Preliminaries
Let $\mathbf{x}\in\mathbb{R}^{d}$ denote an activation vector drawn from a pretrained transformer. A conventional $\operatorname{TopK}$ SAE (Gao et al., 2025) produces a reconstruction $\hat{\mathbf{x}}$ of $\mathbf{x}$ via
@@ 3.1 Preliminaries
where $W_{\text{enc}}\in\mathbb{R}^{F\times d}$ and $W_{\text{dec}}\in\mathbb{R}^{d\times F}$ are full-rank matrices and $\mathbf{f}\in\mathbb{R}^{F}$ retains the $k$ largest activations. The encoder cost therefore scales as $\mathcal{O}(Fd)$ per token.
3.2 KronSAE
In contrast to prior work that (i) accelerates decoding or (ii) sparsifies computation through routing, KronSAE directly reduces the encoder’s computational complexity by decomposing the $W_{\text{enc}}$ matrix into head-wise Kronecker factors and introducing a multiplicative $\operatorname{mAND}$ kernel. We decompose the latent space into $h$ independent components (e.g. heads). Each head $k$ is parameterised by the composition of two thin matrices $P^{k}\in\mathbb{R}^{m\times d}$ (composition base) and $Q^{k}\in\mathbb{R}^{n\times d}$ (composition extension), with dims $m<n\ll d$ and $F=h\,m\,n$. The pre-latents
$\mathbf{p}^{k}=\operatorname{ReLU}(P^{k}\mathbf{x}),\qquad\mathbf{q}^{k}=\operatorname{ReLU}(Q^{k}\mathbf{x})$ (2)
are combined through an element-wise interaction kernel independently in each head,
$\mathbf{z}^{k}_{i,j}:=\operatorname{mAND}(p^{k}_{i},q^{k}_{j}):=\begin{cases}\sqrt{p^{k}_{i}\,q^{k}_{j}},&p^{k}_{i}>0\text{ and }q^{k}_{j}>0,\\ 0,&\text{otherwise},\end{cases}$ (3)
where $\mathbf{z}^{k}\in\mathbb{R}^{m\times n}$. Flattening and concatenating across heads yields post-latents $\mathbf{f}\in\mathbb{R}^{F}$, to which the usual $\operatorname{TopK}$ operator is applied. The $\operatorname{mAND}$ kernel smoothly approximates Boolean AND gate, ensuring non-zero activation only if both inputs are positive, while preserving gradient flow and activation magnitude for stable reconstruction. Further insights into the behavior of $\operatorname{mAND}$ are provided in Appendix C.
The encoder cost per token drops from $\mathcal{O}(Fd)$ to $\mathcal{O}\bigl{(}h(m+n)d\bigr{)}$ (see Appendix A.2). KronSAE thus reduces FLOPs and parameter count without routing overhead, and is orthogonal to existing sparse decoder kernels (Gao et al., 2025) and thus can be combined with them for end-to-end speed-ups.
@@ 4 Experiments
Dictionary size m n h Activation Explained Variance
32768 2 4 4096 $\operatorname{mAND}(u,v)$ $\mathbf{0.83356}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.82669
$u\cdot v$ 0.82373
4 8 1024 $\operatorname{mAND}(u,v)$ $\mathbf{0.82197}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.81912
$u\cdot v$ 0.81432
65536 2 4 8192 $\operatorname{mAND}(u,v)$ $\mathbf{0.84445}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.83276
$u\cdot v$ 0.82974
4 8 2048 $\operatorname{mAND}(u,v)$ $\mathbf{0.83503}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.82968
$u\cdot v$ 0.8251
Table 1: Explained variance for KronSAE using different composition activations under a 1B-token training budget. Bold values indicate the best performance of $\operatorname{mAND}$ across composition kernels.
We train SAEs on the residual streams of Qwen-2.5-1.5B-Base (Yang et al., 2024), Pythia-1.4B-deduped (Biderman et al., 2023), and Gemma-2-2B (Team et al., 2024) language models. Activations are collected on FineWeb-Edu, a filtered subset of educational web pages from the FineWeb corpus (Penedo et al., 2024). We measure reconstruction quality via explained variance (EV),
@@
so that $1.0$ is optimal, and use automated interpretability pipeline (Bills et al., 2023; Paulo et al., 2024) and SAE Bench (Karvonen et al., 2025) to evaluate properties of SAE features. Our experiments (see detailed setup in Appendices A and D) address three questions:
@@
    Does KronSAE yield higher EV than a TopK SAE under fixed compute?
@@
    How do these choices affect properties and interpretability of learned features?
4.1 Scaling Laws
We examine the scaling behavior of KronSAE against a conventional TopK SAE on Qwen-1.5B under equivalent FLOPs budgets. Models are grouped by dictionary size $F\in\{2^{15},2^{16},2^{17}\}$, and each KronSAE configuration $(m,n,h)$ matching $F$ hidden features is trained under token budgets adjusted to match the per-step FLOPs of training a TopK SAE for 100 M, 500 M, and 1000 M tokens (see FLOPs calculation details in Appendix A).
Figure 1 plots explained variance versus trainable parameter count across these isoflops budgets. Under a 100 M token budget, all KronSAE variants exceed the TopK baseline in reconstruction quality, achieving up to a 4.3% increase in explained variance with roughly 54.7% fewer parameters. At 500 M tokens, KronSAE maintains  0.8% advantage and 43.8% fewer parameters across most model sizes, and the compositional head structure keeps smaller configurations on or above the baseline Pareto front. With 1000 M tokens, although the TopK SAE narrows the gap, KronSAE still matches baseline reconstruction quality with approximately 46.1% fewer parameters. Smaller composition base dimension ($m=2$) show especially strong returns under tight compute constraints. We observed similar results on Pythia-1.4B (see section A.3).
@@ 4.2 Ablations
Figure 2: Iso-FLOPs ablation: EV against head count $h$ and base dimension $m$ for KronSAE under 100M, 500M, and 1B token budgets. Higher $h$ and smaller $m$ yield improved reconstruction quality.
@@
To isolate the impact of our $\operatorname{mAND}$ operator, we compare it to two simpler interaction kernels: (i) the element-wise product of ReLUs, $\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$, and (ii) the raw product $u\cdot v$. As reported in Table 1, under a 1B-token training budget and for both $(m,n,h)=(2,4,4096)$ and $(4,8,1024)$ configurations, the $\operatorname{mAND}$ variant consistently achieves the highest explained variance, outperforming the alternatives by a clear margin.
Decomposition hyperparameters.
Using the iso-FLOPs setup from Section 4.1, we systematically vary the number of heads $h$ and the per-head base dimension $m$ (with $n=F/(mh)$). Figure 2 shows that for large training budgets (500M–1B tokens), smaller $m$ (and thus larger $n$) yields higher reconstruction quality, since a smaller base dimension frees capacity for more expressive extension features. Under the tighter 100M-token regime, however, configurations with $m=4$ outperform those with $m=2$ or $m=8$, indicating a trade-off between per-head representational richness and data efficiency. Furthermore, fixing $m$ and increasing $h$ yields near-linear gains in explained variance. This corresponds to our intuition that each head acts as a component of the feature-space decomposition, efficiently storing a group of correlated features and allowing more fine-grained routing of activation patterns.
Sparsity and depth-wise position.
Finally, we assess KronSAE’s robustness to different sparsity levels and layer depths. Figure 3(a) compares KronSAE and TopK SAE across a range of $L_{0}$ sparsity settings on the 14th layer of Qwen-2.5-1.5B and the 12th layer of Gemma-2-2B; Figure 3(b) evaluates performance across layers in Qwen-2.5-1.5B. In every case, KronSAE matches or exceeds the reconstruction quality of the TopK baseline under the same FLOPs budget, demonstrating that our Kronecker-factorized encoder maintains its advantages regardless of sparsity level or depth.

### v3

# Kronecker Factorization Improves Efficiency and Interpretability of Sparse Autoencoders

@@ Abstract
Sparse Autoencoders (SAEs) have demonstrated significant promise in interpreting the hidden states of language models by decomposing them into interpretable latent directions. However, training and interpreting SAEs at scale remains challenging, especially when large dictionary sizes are used. While decoders can leverage sparse-aware kernels for efficiency, encoders still require computationally intensive linear operations with large output dimensions. To address this, we propose KronSAE – a novel architecture that factorizes the latent representation via Kronecker product decomposition, drastically reducing memory and computational overhead. Furthermore, we introduce mAND, a differentiable activation function approximating the binary AND operation, which improves interpretability and performance in our factorized framework.
@@ 1 Introduction
Interpreting large language models and their embeddings in particular remains a central challenge for transparency and controllability in AI systems (Elhage2023; Heap2025). Sparse autoencoders (SAEs) have emerged as powerful tools for uncovering human‐interpretable features within neural activations by enforcing activation sparsity to induce discrete‐style dictionaries (Elhage2023; gao2025scaling; Cunningham2024). These dictionaries facilitate circuit‐level semantic analysis (marks2025sparse) and concept discovery, enabling fine‐grained probing of model internals (Elhage2023; Cunningham2024).
However, naively scaling SAEs to the widths demanded by modern transformers leads to prohibitive compute costs, limiting their applicability to large-scale interpretation experiments. Gated SAEs (DBLP:conf/nips/RajamanoharanCS24) address this by learning continuous sparsity masks via lightweight gating networks, and Switch SAEs (mudide2025efficient) leverage conditional computation by routing activations among smaller expert SAEs, reducing computation by activating only a subset of experts per input. gao2025scaling propose an acceleration of TopK SAE that utilizes an optimized kernel based on efficient sparse–dense matrix multiplication. Encoder remains unoptimized: it still performs a dense projection into the full dictionary, incurring high computational cost and limiting scalability.
Another limitation of SAEs is the absence of structure within learned latents in its classical design. While mudide2025efficient address this via expert subnetworks, bussmann2025learning imposes feature hierarchy via nested dictionaries structure and improves the interpretability of SAE latents.
In this paper we address both these directions and introduce KronSAE, an encoder that is applicable to many existing SAE architectures. By decomposing the latent space into head‐wise Kronecker factors and including differentiable logical AND‐like gating mechanism, KronSAE reduces both parameters and compute overhead while preserving reconstruction fidelity and improves interpretability.
@@
    We identify the encoder projection as one of the principal scalability bottlenecks in sparse autoencoders, demonstrating that targeted encoder optimizations can significantly improve computational performance while maintaining reconstruction quality.
@@
    We propose KronSAE, a Kronecker-factorised sparse autoencoder equipped with the novel $\operatorname{mAND}$ activation function. The design reduces encoder cost and is compatible with existing sparse decoder kernels, loss function designs and mixture of experts architecture.
@@
    We show on multiple language models that KronSAE decreases feature absorption and yields more interpretable latents under fixed compute.
@@ 2 Related Work
Early work demonstrated that SAEs can uncover human-interpretable directions in deep models (Cunningham2024; templeton2024scaling), but extending them to the large latent sizes ($F\gg d$) for modern language models with large hidden dimension $d$ is expensive: each forward pass still requires a dense $\mathcal{O}(F\,d)$ encoder projection. Most prior optimization efforts focus on the decoder side: for example, gao2025scaling introduce a fused sparse–dense TopK kernel that reduces wall-clock time and memory traffic, while DBLP:conf/nips/RajamanoharanCS24 decouple activation selection from magnitude prediction to improve the $\ell_{0}$–MSE trade-off. Separately, rajamanoharan2024jumpingahead propose JumpReLU, a nonlinearity designed to mitigate shrinkage of large activations in sparse decoders.
@@
These schemes avoid instantiating a full dictionary per token by routing inputs to a subset of expert SAEs via a lightweight gating network (mudide2025efficient) by employing the Mixture-of-Experts ideas (Shazeer2017), but still incur a dense per-expert encoder projection, leaving the encoder as the primary bottleneck.
Factorizations and Logical Activation Functions.
Tensor product representations (smolensky1990tensorproductrepresentations) have been utilized to represent the compositional structures in dense embeddings, closely resembling our idea of AND-like compositions, and recent studies have been extended TPR to study the transformer hidden states and in-context learning  (soulos-etal-2020-discovering; smolensky2024mechanismssymbolprocessingincontext). Separately, tensor-factorization methods have been used to compress large weight matrices in language models (edalati2021kroneckerdecompositiongptcompression; wang2023learning), and structured matrices have been utilized to improve training efficiency and might be used to impose the inductive bias (dao2022monarch). In parallel, differentiable logic activations were introduced to approximate Boolean operators in a smooth manner (Lowe2021). Our method synthesizes these lines of work: we embed a differentiable AND-like gate into a Kronecker-factorized efficient encoder to build compositional features while preserving end-to-end differentiability.
@@ 3 Method
Preliminaries.
Let $\mathbf{x}\in\mathbb{R}^{d}$ denote an activation vector drawn from a pretrained transformer. A conventional $\operatorname{TopK}$ SAE (gao2025scaling) produces a reconstruction $\hat{\mathbf{x}}$ of $\mathbf{x}$ via
@@
where $W_{\text{enc}}\in\mathbb{R}^{F\times d}$ and $W_{\text{dec}}\in\mathbb{R}^{d\times F}$ are dense matrices and $\mathbf{f}\in\mathbb{R}^{F}$ is a sparse vector retaining only the $K$ largest activations. The encoder cost therefore scales as $\mathcal{O}(Fd)$ per token.
KronSAE.
Our method reduces the encoder’s computational cost while also enforcing compositional structure of the latents. We decompose the latent space into $h$ independent heads, and each head $k$ is parameterised by the composition of two thin matrices $P^{k}\in\mathbb{R}^{m\times d}$ (composition base) and $Q^{k}\in\mathbb{R}^{n\times d}$ (composition extension), with dimensions $m\leq n\ll d$ and F=hmnF=h\,m\,n. The pre‐latents
$\mathbf{p}^{k}=\operatorname{ReLU}(\mathbf{u}^{k})\qquad\text{and}\qquad\mathbf{q}^{k}=\operatorname{ReLU}(\mathbf{v}^{k}),$ (2)
with $\mathbf{u}^{k}=P^{k}\mathbf{x}$ and $\mathbf{v}^{k}=Q^{k}\mathbf{x}$ acting as the elements from which compositional features would be built, are combined through an element‐wise interaction kernel independently in each head:
$z^{k}_{i,j}:=\operatorname{mAND}(u^{k}_{i},v^{k}_{j}):=\begin{cases}\sqrt{u^{k}_{i}\,v^{k}_{j}},&u^{k}_{i}>0\text{ and }v^{k}_{j}>0,\\ 0,&\text{otherwise},\end{cases}$ (3)
where $\mathbf{z}^{k}\in\mathbb{R}^{m\times n}$; it is then flattened in a row-major order to a vector equivalent to the element-wise square root of the Kronecker product of $\mathbf{p}^{k}$ and $\mathbf{q}^{k}$. Assuming that $\operatorname{vec}(\cdot)$ is in row-order, we get
$\operatorname{vec}(\mathbf{z}^{k})=\sqrt{\operatorname{vec}\left(\mathbf{p}^{k}(\mathbf{q}^{k})^{\top}\right)}=\sqrt{\mathbf{p}^{k}\otimes\mathbf{q}^{k}}$ (4)
Concatenating heads and applying $\operatorname{TopK}$ yields post‐latents $\mathbf{f}\in\mathbb{R}^{F}$. By ensuring active $z_{i,j}$ only when $p_{i}>0$ and $q_{j}>0$ we directly force the AND-like interaction: let $\mathcal{P}$ and $\mathcal{Q}$ be the sets of hidden state vectors on which $p_{i}$ and $q_{j}$, respectively, are active, and let $\mathcal{F}$ be the set of inputs on which $z_{i,j}>0$, then it is true that $\mathcal{P}\cap\mathcal{Q}=\mathcal{F}$. The square root in equation 3 prevents activation value explosion when both pre-latents are active. See additional discussion in Appendix C.
The encoder cost per token drops from $\mathcal{O}(Fd)$ to $\mathcal{O}\bigl(h(m+n)d\bigr)$ (see Appendix A.2). KronSAE thus reduces FLOPs and parameter count without routing overhead, and is orthogonal to existing sparse decoder kernels (gao2025scaling) and thus can be combined with them for end-to-end speed-ups.
@@ 4 Experiments
We train SAEs on the residual streams of Qwen-2.5-1.5B-Base (Yang2024), Pythia-1.4B-deduped (biderman2023pythiasuiteanalyzinglarge), and Gemma-2-2B (gemmateam2024gemma2improvingopen) language models. Activations are collected on FineWeb-Edu, a filtered subset of educational web pages from the FineWeb corpus (penedo2024finewebdatasetsdecantingweb). We measure reconstruction quality via explained variance (EV),
@@
so that $1.0$ is optimal, and use automated interpretability pipeline (bills2023llmcanexplain; paulo2024automatically) and SAE Bench (karvonen2025saebench) to evaluate properties of SAE features. We aim for the needs of resource-constrained mechanistic interpretability research where efficiency and interpretability are in favor rather than top reconstruction performance, and 100M-820M token budgets are widely adopted (bussmann2025learning; kharlapenko2025scaling; Heap2025; mudide2025efficient; karvonen2025saebench), so we choose 125M, 500M, 1B and 2B token budgets for the experiments.
Our experiments (see detailed setup in Appendices A and D) address three questions:
@@
    Does KronSAE maintain EV comparable to baseline SAEs under fixed compute?
@@
    How do these choices affect properties and interpretability of learned latents?
@@ 4.1 Ablations
We employ the iso-FLOPs setup: for each KronSAE variant of dictionary size $F$ we allocate the same amount of FLOPs as was spent for the training of TopK SAE for token budget $T$ and same $F$.
Reconstruction performance.
As indicated on Figure 1, KronSAE achieves on-par performance with TopK given lower number of trainable parameters and outperforms Matryoshka SAE. The performance gap narrows when increasing the dictionary size, which indicate the potential scalability of our method for large dictionaries. See also result on Gemma-2 2B model in Appendix A.3.
Figure 1: Maximum performance for KronSAE vs. TopK SAE vs. Matryoshka TopK SAE on Qwen-1.5B for different dictionary sizes $F$ and budgets in iso-FLOP setting. KronSAE with lower number of parameters is on-par with the baseline, and the gap narrows with larger dictionary size.
Decomposition hyperparameters.
We systematically vary the number of heads $h$ and the per‐head base dimension $m$ (with $n=F/(mh)$) under the iso-FLOPs setup. From the Figure 2, we conclude that lower $m$ and higher $h$ consistently yields higher reconstruction quality, due to flexibility of pre-latents - as we show in section 5.3, they must be either expressive or fine-grained enough (low $m$, $n$ and high $h$) to efficiently represent the semantics. See also Appendices A.5 and A.8.
Figure 2: Dependency of EV on head count $h$ (on the x-axis) and base dimension $m$ under 500M and 1B token budgets in iso-FLOPs setup. Higher $h$ and smaller $m$ yield improved reconstruction quality because of higher expressivity of pre-latents to encode semantics and increasing trainable parameters.
@@
To isolate the impact of our $\operatorname{mAND}$ operator, we compare it to two simpler interaction kernels: (i) the element‐wise product of ReLUs, $\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$, and (ii) the raw product $u\cdot v$. As reported in Table 1, under a 125M training budget, the $\operatorname{mAND}$ variant achieves the highest explained variance. More description of the $\operatorname{mAND}$ is provided in the Appendix C. See also our experiments where we replace TopK with JumpReLU in Appendix A.6.
Dictionary size m n h Activation Explained Variance
32768 2 4 4096 $\operatorname{mAND}(u,v)$ $\mathbf{0.8336}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.8267
$u\cdot v$ 0.8237
4 8 1024 $\operatorname{mAND}(u,v)$ $\mathbf{0.8220}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.8191
$u\cdot v$ 0.8143
65536 2 4 8192 $\operatorname{mAND}(u,v)$ $\mathbf{0.8445}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.8328
$u\cdot v$ 0.8297
4 8 2048 $\operatorname{mAND}(u,v)$ $\mathbf{0.8350}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.8297
$u\cdot v$ 0.8251
Table 1: Performance of different composition activations under a budget of 125M tokens.
Sparsity Analysis.
To evaluate performance across different sparsity budgets ($\ell_{0}=\{16,32,64,128\}$), we compare multiple SAE baselines (TopK SAE, Matryoshka SAE, Switch SAE) against their KronSAE variants at dictionary size $F=2^{16}$ trained on 500 million tokens, under an iso-FLOPs budget matched to the TopK baseline. As shown in Figure 3, KronSAE achieves comparable reconstruction fidelity (measured by explained variance) to all baseline methods across all sparsity levels.Results for smaller dictionary size $F=2^{15}$ are presented in Appendix A.4.
Figure 3: Maximum performance for baselines and their KronSAE modifications for different sparsity levels in iso-FLOP setting. KronSAE variants, despite using fewer trainable parameters, achieve reconstruction quality comparable to or better than the unmodified baselines.
Layerwise performance.
Additionally, we evaluate performance across different layers in Qwen-2.5-1.5B. In every case, KronSAE matches the reconstruction quality of the TopK baseline, demonstrating that our Kronecker‐factorized encoder maintains its performance regardless of depth. This setup and corresponding results are described in greater detail in Appendix A.7.